Fermi National Laboratory
 

Accelerator Update for June 26, 2006 - June 27, 2006

 

Monday June 26

The day shift began with Operations monitoring store 4798 and a Recycler stash, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, with the Antiproton Source (Pbar) stacking, with the Linac, Booster, Main Injector (MI), Switchyard (SY), and Recycler operating normally, and with MiniBooNE and NuMI taking beam.

Operations terminated store 4798 at 9:28 AM.

At 9:44 AM, Operations halted MI beam so an expert could repair a RF station (MIRF13). The expert completed the repair at 10:05 AM.

Operation put the Tevatron (TeV) into shot setup at 10:05 AM.

Operations established store 4799 around 12:20 PM with an initial luminosity of 106.34E30. Pbar resumed stacking.

Operations established beam to the Meson MTest experiment at 12:40 PM.

At 12:45 PM, a Linac RF station (KFR1) tripped off twice in ten minutes and then had a problem with its phase wandering.

Recycler began an energy matching study at 12:57 PM.

At 1:21 PM, KRF1 tripped and wouldn’t reset. Linac technicians adjusted its voltage reference and it resumed normal operations.

Operations began preparing the Recycler at 2:14 PM to stash antiprotons.

At 2:39 PM, Linac experts reported that they needed to change a current monitor for KRF1.

At 2:59 PM, the first attempt to stash antiprotons failed; they never got out of Pbar. Operators investigated.

Since the stash was on hold, Operations let the Linac techs begin work on KRF1 at 3:01 PM. This work held off all beam.

Operations reported 4 PM that they found and corrected a voltage setting for a corrector magnet power supply (VT906), which had hindered the Recycler stash.

The stash was ended and Pbar resumed stacking at 4:03 PM.

NuMI and MiniBooNE resumed taking beam at 4:05 PM.

MTest resumed taking beam at 4:10 PM.

At 4:18 PM, Operations reported a vacuum problem in Booster. Experts replaced a vacuum pump power supply. Beam resumed at 4:37 PM.

At 8:47 PM, a person reported a nitrogen leak from a Dewar outside of the sector A0 service building. A cryo system tech tried to fix the leak, but couldn’t. He valved in a different Dewar. The leaking Dewar will be drained and repaired on Tuesday.

The Recycler began preparing to stash at 9:31 PM. The stash ended and Pbar and MTest resumed normal operations at 10:55 PM.

Tuesday June 27

The midnight shift began with Operations monitoring store 4799 and the stash, with Pbar stacking, with normal Linac, MI, Recycler, SY, and Booster operations, and with MTest, MiniBooNE, and NuMI taking beam.

MTest experimenters reported at 12:59 AM that they were finished with beam for the night.

The Recycler began preparing to stash at 4:01 AM. The stash ended and Pbar resumed stacking at 5:50 AM.

Operations established beam to MTest at 7:40 AM.

The Plans for Tuesday

The plans for today are to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/27/2006    email Fermilab